Can you show me step by step how to solve 4-ln(5-x)=5/2

Answer by lwsshak3(11628)

Can you show me step by step how to solve
4-ln(5-x)=5/2
ln(5-x)=4-5/2=8/2-5/2=3/2
ln(5-x)=3/2
convert to exponential form:
e^(3/2)=5-x
x=5-e^(3/2)
x≈0.5183